# Susceptibility for extremely low external fluctuations and critical behaviour of Greenberg-Hastings neuronal model

### v1

@@ I Introduction
The Greenberg-Hastings (GH) cellular automaton is one of the simplest neural network models displaying cooperative emergent properties [1]. Despite its simplicity, it exhibits a rich dynamical behaviour, at least as far as criticality is concerned, including network topology dependent first and second order phase transitions [2]-[5], as well as tricritical behavior when both inhibitory and excitatory units are considered [6]. A central quantity in all these phenomena is the activation threshold $T$ of the artificial neurons, which acts as a control parameter for all the phase transitions the system undergoes for certain network topologies (in the case of purely excitatory units, it is the only control parameter). Then, two different (although closely related) quantities exhibit the typical behaviour of an order parameter as $T$ changes: the average fraction of neurons in the largest cluster of excited neurons (thus describing a percolation-like phase transition [1, 2, 4]) and the average total fraction of excited neurons [3, 4]. Both quantities become negligible above a certain characteristic (i.e., critical) value of $T$ and take large values below it. The curious fact is that, at least for some network topologies, both phenomena seem to happen at different critical values of the control parameter $T$, even in the infinite size limit [4]. Moreover, the susceptibilities associated with both order parameters exhibit different scaling properties: while the average cluster size (the standard susceptibility in percolation phenomena) exhibits a size-dependent peak with associated critical exponents in the standard mean field percolation universality class [2], the fluctuation susceptibility associated with the total activity seems to exhibit a size-independent peak [4]. Understanding the last fact is expected to give an important clue to the general comprehension of this complex critical scenario. The main part of the present work focuses on the finite size susceptibility problem.
@@ II The Model
[…]
@@
×$\times\quantity[1-\Theta\quantity(\sum_{j}W_{ij}\delta(x_{j},1)-T)]$
$P_{i}(1\rightarrow 2)$ $=$ 11 11
@@
where $\Theta$ is the Heaviside function and $\delta(x,y)$ is the Kroenecker delta function. $W_{ij}$ is a symmetric synaptic matrix, and the sum is carried out over the nearest neighbours of the node $i$. Notice that $P_{i}(0\rightarrow 1)$ is written as $1$ minus the probability of not being activated, which is the product of probabilities of not having a spontaneous activation, $(1-r_{1})$, and not having an activation due to active neighbours. Attempting to mimic the weight distribution of the human connectome [12], synaptic weights are quenched variables chosen from an exponential distribution
@@
[…]
@@
In a second-order phase transition, this quantity is expected to present a size-dependent maximum which scales as $\max\chi\sim N^{\gamma^{\prime}/\nu d}$ for $N\gg 1$, where $N$ is the number of nodes in the network, $\gamma^{\prime}$ and $\nu$ are the critical exponents associated with the infinite size divergence of $\chi$ and the correlation length respectively and $d$ is the effective dimension of the model [15]. Finally, the first coefficient of the autocorrelation function,
$AC(1)=\frac{\langle a(t+1)-\langle a\rangle\rangle.\langle a(t)-\langle a\rangle\rangle}{Var(a)},$ (5)
@@ III Absorbing critical behaviour without spontaneous activity
Figure 1: Susceptibility as a function of the threshold $T$ for different system sizes $N$ when $r_{1}=0$, obtained from different quasi-stationary algorithms. The simulations correspond to the GH model defined on a Watts-Strogatz network of size $N$ with average degree $\langle k\rangle=12$ and rewiring probability $\pi=0.6$. The continuous lines correspond to fittings using appropriate fitting functions. The insets show the maxima of the different curves as a function of $N$ estimated from the fittings, with power law fittings (red lines) $\max\chi\sim N^{\gamma^{\prime}/\nu d}$. The critical exponents estimations are shown in Table 1. (a) Curves obtained using the reactivation algorithm. (b) Curves obtained using the fixed time algorithm.
@@
Figure 2: (a) Pseudo critical threshold $T_{c}^{*}$ estimated from the fittings of Fig.1 and (b) activity $f_{a}$ evaluated at $T_{c}^{*}$, as a function of network size $N$ for $r_{1}=0$. Blue circles correspond to data from the reactivation algorithm and red squares to the fixed time algorithm. The continuous lines are fitting curves using $T^{*}_{c}\sim T_{c}-N^{-1/\nu d}$ and $f_{a}(T=T^{*}_{c})\sim N^{-\beta/\nu d}$. The estimated values of $T_{c}$, $1/\nu d$ and $\beta/\nu d$ from the fittings are shown in Table 1. Inset: $T_{c}-T_{c}^{*}$ in log-log scale. Dashed lines are the corresponding power law fittings.
@@
We run simulations of the GH model defined on a Watts-Strogatz network of size $N$ with average degree $\langle k\rangle=12$ and rewiring probability $\pi=0.6$, from which we obtained both the average network activity $f_{a}$ and the susceptibility $\chi$ from Eq.(4) using both algorithms for $r_{1}=0$. The obtained susceptibility as a function of the threshold $T$ for different system sizes is shown in Fig.1. We verified the presence of a size divergent peak of $\chi$. All the curves were fitted using appropriate fitting functions, from which the maxima of the curves were estimated as well as their location (pseudo critical points) $T_{c}^{*}(N)$. Using these values, we calculated the order parameter (activity) at the pseudo critical point $f_{a}(T=T_{c}^{*})$. The results for $T_{c}^{*}$ and $f_{a}(T=T_{c}^{*})$ as a function of $N$ are shown in Fig.2. We verified the expected finite size scaling behaviours in a critical phenomenon [15]: $\max\chi\sim N^{\gamma^{\prime}/\nu d}$, $T^{*}_{c}\sim T_{c}-N^{-1/\nu d}$ and $f_{a}(T=T^{*}_{c})\sim N^{-\beta/\nu d}$. The estimated values of $T_{c}$ and the critical exponents are summarised in Table 1.
@@
Finally, the third method consists of running standard simulations for decreasing values of $r_{1}$, considering the $r_{1}\to 0$ limit. In Fig.3 we illustrate the typical behaviour of $\chi$ as a function of $T$ for different system sizes for three typical values of $r_{1}$. While for relatively large values of $r_{1}$ we recover the absence of scaling [4], as $r_{1}$ decreases, we see a clear emergence of a divergent peak with $N$. Following the same procedure as with the previous two methods, we estimated the critical exponents $\gamma^{\prime}/\nu d$, $\beta/\nu d$ and $1/\nu d$, for different values of $r_{1}$. The results are shown in Fig.4. We have a clear convergence when $r_{1}\to 0$ of all the exponents to values that are similar to the ones obtained by the two $r_{1}=0$ algorithms. The average values when $r_{1}\to 0$ are compared with those estimations in Table 1. We see that the three estimations are consistent, giving approximate values $\beta\approx 0.3$ and $\gamma^{\prime}\approx 0.25$ which, to the best of our knowledge, do not agree with any previously known universality class. Notice that in all the cases the expected scaling law [15]
$\frac{\gamma^{\prime}}{\nu d}+2\frac{\beta}{\nu d}=1$ (6)
is approximately satisfied (see Table 1).
@@
Figure 3: Susceptibility as function of the threshold $T$ in Watts-Strogatz networks with $\langle k\rangle=12$ and $\pi=0.6$ for different network sizes $N$ and three typical values of the spontaneous activation probability: (a) $r_{1}=10^{-3}$, (b) $r_{1}=5\times 10^{-4}$ and (c) $r_{1}=10^{-5}$. The continuous lines are a guide to the eye.
@@ IV Activation mechanisms
Assuming that each neuron interacts with the average activity on the network, when $T$ is high enough, we can approximate $1-p_{n}(t)\simeq[1-a(t)P(W>T)]^{k}$, where $1-a(t)P(W>T)$ is the probability of not being activated by a single active neighbour, and $k$ is the number of neighbour neurons (i.e., we do not consider here the possibility of a cooperative activation).
@@
Similarly, the equation accounts for random spontaneous activation and also for activation due to a single active presynaptic neuron (single activation, $a^{(s)}$). Cooperative activation due to the contribution of several presynaptic contributions $a^{(c)}$ is not considered there. In that sense, notice that the equation is quite similar to that derived in Ref.[7] for a different model which does not include cooperative activation at all. It is easy to see that Eq.(8) predicts a critical point at $r_{1}=0$ and $T_{c}=\ln(k)/\lambda$ where, close to the transition, we have $f_{a}(T,r_{1}=0)\sim(T-T_{c})^{\beta_{MF}}$, $\chi^{\prime}(T,r_{1}=0)={\partial f_{a}\over\partial r_{1}}|_{r_{1}=0}\sim(T-T_{c})^{-\gamma_{MF}}$ and $f_{a}(T=T_{c},r_{1})\sim r_{1}^{1/\delta_{MF}}$. The critical exponents are those of mean field directed percolation exponents: $\beta_{MF}=1$, $\gamma_{MF}=1$ and $\delta_{MF}=2$, in agreement with the results of Ref.[7]. We also see that, consistently with the results of section III, the spontaneous activation probability $r_{1}$ behaves as an external field conjugated to the order parameter $f_{a}$ of the continuous phase transition.
@@
Figure 7: Activities, susceptibility and $AC(1)$ as a function of $T$ for $r_{1}=10^{-5}$, $N=10000$, $\pi=0.6$ and different values of $\langle k\rangle$. Each column corresponds to one value of $\langle k\rangle$. From left to right $\langle k\rangle=2,8,12,40$. Upper row shows the activities as a function of $T$. Intermediate rows show the susceptibility $\chi$ and the first autocorrelation coefficient $AC(1)$. Blue dashed lines mark the pseudo critical thresholds $T_{c}^{*}$ obtained from $AC(1)$ peaks. The colour curves in the inset correspond to larger system sizes ($N=80000$ to $1280000$) and show that the scaling peak of $\chi$ for $\langle k\rangle=8$ only emerges for large enough network sizes. The lower row shows the ratio among cooperative and single contributions to the activity, $f_{a}^{c}/f_{a}^{s}$ (spontaneous activation $f^{(sp)}=f_{a}-f_{a}^{(c)}-f_{a}^{(s)}$ is negligibly small and consequently not plotted). All simulations were averaged over at least 30 networks.
@@
In Fig.8 we compare the estimation of $T_{c}$ from the maximum of $AC(1)$ and the non- trivial (i.e., that scales with $N$) peak of the susceptibility as a function of $\langle k\rangle$ for $N=10000$ and $r_{1}=0$, using the reactivation algorithm (see section III). Checks were made for some values of $\langle k\rangle$ by extrapolating the $N\to\infty$ limit of $T_{c}^{*}$ in the continuous regime and also with standard simulations with $N=10000$ and $r_{1}=10^{-5}$. The differences were negligible in all the cases. We see that the estimations obtained from both quantities ($AC(1)$ and $\chi$) are indistinguishable for all values of $\langle k\rangle$. However the peak of $AC(1)$ is broad (i.e., it decays slowly as $|T-T_{c}|$ is increased) while the peak of $\chi$ may be narrow for some network topologies. Because of that, a reasonable estimate of $T_{c}$ can be found from $AC(1)$, for all studied values of $k$, even for small values of $N$.

### v2

@@ I Introduction
The Greenberg-Hastings (GH) cellular automaton is one of the simplest neural network models displaying cooperative emergent properties [1]. Adopting the G&H formalism has been well received to model large-scale brain dynamics since the work of Haimovici et al [1] a decade ago, which successfully replicated several key observations in human brain imaging. Its simplicity allowed it to focus the attention on the role of the critical dynamics in reproducing the large-scale organization of the spatio-temporal patterns observed in the human brain under resting conditions. Since then, scores of papers have used the same approach to study the brain dynamics under different conditions as sleep, anaesthesia, coma, stroke, influence of psychedelics, etc.. Despite its simplicity, the GH neural model exhibits a rich dynamical behaviour, at least as far as criticality is concerned, including network topology dependent first and second order phase transitions [2]-[5], as well as tricritical behaviour when both inhibitory and excitatory units are considered [6]. A central quantity in all these phenomena is the activation threshold $T$ of the artificial neurons, which acts as a control parameter for all the phase transitions the system undergoes for certain network topologies (in the case of purely excitatory units, it is the only control parameter). Then, two different (although closely related) quantities exhibit the typical behaviour of an order parameter as $T$ changes: the average fraction of neurons in the largest cluster of excited neurons (thus describing a percolation-like phase transition [1, 2, 4]) and the average total fraction of excited neurons [3, 4]. Both quantities become negligible above a certain characteristic (i.e., critical) value of $T$ and take large values below it. The curious fact is that, at least for some network topologies, both phenomena seem to happen at different critical values of the control parameter $T$, even in the infinite size limit [4]. Moreover, the susceptibilities associated with both order parameters appear to exhibit different scaling properties: while the average cluster size (the standard susceptibility in percolation phenomena) exhibits a size-dependent peak with associated critical exponents in the standard mean field percolation universality class [2], the fluctuation susceptibility associated with the total activity seems to exhibit a size-independent peak [4]. Understanding the last fact is expected to give an important clue to the general comprehension of this complex critical scenario. The main part of the present work focuses on the finite size susceptibility problem.
@@ II The Model
We considered a neural model first introduced by Haimovici et al. [1] on a network derived from the human connectome [12] and later implemented in a small-world network by Zarepour et al. [2]. The dynamics is that of a Greenberg-Hastings cellular automaton [13], where each site $i$ of a network has associated a three-state variable $x_{i}=0,\,1,\,2$, representing the following dynamical states: quiescent ($x_{i}=0$), excited ($x_{i}=1$) and refractory ($x_{i}=2$). The network is constructed from a one-dimensional ring with homogeneous degree $k$. Following the usual Watts-Strogatz (WS) recipe [14], bonds are rewired with probability $\pi$, so the average degree $\langle k\rangle$ remains constant. The model follows a parallel dynamics in a discrete-time $t$ as follows: (a) each quiescent neuron can become excited with a small spontaneous activation probability $r_{1}$, or if the sum of their active neighbours, weighted through the connectivity matrix, surpasses an activation threshold $T$ (the control parameter of this model), which is the same for all neurons; (b) active neurons become refractory after one step, and (c) refractory neurons become quiescent with probability $r_{2}$ per unit time (so, the average refractory period has a duration of $1/r_{2}$). The transition probabilities for the i-th site are given by
@@
$\times\quantity[1-\Theta\quantity(\sum_{j}W_{ij}\delta(x_{j},1)-T)]$
$P_{i}(1\rightarrow 2)$ $=$ $1$
@@
where $\Theta$ is the Heaviside function and $\delta(x,y)$ is the Kronecker delta function. $W_{ij}$ is a symmetric synaptic matrix, and the sum is carried out over the nearest neighbours of the node $i$. Notice that $P_{i}(0\rightarrow 1)$ is written as $1$ minus the probability of not being activated, which is the product of probabilities of not having a spontaneous activation, $(1-r_{1})$, and not having an activation due to active neighbours. Attempting to mimic the weight distribution of the human connectome [12], synaptic weights are quenched variables chosen from an exponential distribution
@@
is the instantaneous fraction of excited neurons, and $\langle\cdots\rangle$ stands for both a temporal average and over the quenched disorder. This quantity acts as an order parameter showing high and low activity regimes when the control parameter $T$ is varied. The fluctuation susceptibility of $f_{a}$ is given by
@@
In a second-order phase transition, this quantity is expected to present a size-dependent maximum which scales as $\max\chi\sim N^{\gamma^{\prime}/\nu d}$ for $N\gg 1$, where $N$ is the number of nodes in the network, $\gamma^{\prime}$ and $\nu$ are the critical exponents associated with the infinite size divergence of $\chi$ and the correlation length respectively and $d$ is the effective dimension of the model [15] (namely, $d$ equals the spatial dimension if smaller than the upper critical dimension $d_{c}$ and $d=d_{c}$ otherwise). WS networks do not have a well defined dimension [16]. The spatial dimension $D$, defined as $D=d(\log N)/d(\log l)$, where $l$ is the linear dimension of the network, behaves (for a WS network constructed from a one dimensional ring) as $D\sim\log\left(\pi\langle k\rangle N\right)$ for $\pi\langle k\rangle N\gg 1$ [16]. Therefore, in the thermodynamic limit, they become infinite dimensional. Even in finite size systems, for the range of network sizes of our simulations, we have $D>10$, so we can assume $d=d_{c}$.
Finally, the first coefficient of the autocorrelation function,
$AC(1)=\frac{\langle a(t+1)-\langle a\rangle\rangle\langle a(t)-\langle a\rangle\rangle}{Var(a)},$ (5)
@@ III Absorbing critical behaviour without spontaneous activity
Figure 1: Susceptibility as a function of the threshold $T$ for different system sizes NN when $r_{1}=0$, obtained from different quasi-stationary algorithms. The simulations correspond to the GH model defined on a Watts-Strogatz network of size $N$ with average degree $\langle k\rangle=12$ and rewiring probability $\pi=0.6$. The continuous lines correspond to fittings using appropriate fitting functions (asymmetric double sigmoidal functions). The insets show the maxima of the different curves as a function of NN estimated from the fittings, with power law fittings (red lines) $\max\chi\sim N^{\gamma^{\prime}/\nu d}$. The critical exponents estimations are shown in Table 1. (a) Curves obtained using the reactivation algorithm. (b) Curves obtained using the fixed time algorithm.
@@
Figure 2: (a) Pseudo critical threshold $T_{c}^{*}$ estimated from the fittings of Fig.1 and (b) activity $f_{a}$ evaluated at $T_{c}^{*}$, as a function of network size NN for $r_{1}=0$. Blue circles correspond to data from the reactivation algorithm and red squares to the fixed time algorithm. The continuous lines are fitting curves using $T^{*}_{c}\sim T_{c}-N^{-1/\nu d}$ and $f_{a}(T=T^{*}_{c})\sim N^{-\beta/\nu d}$. The estimated values of $T_{c}$, $1/\nu d$ and $\beta/\nu d$ from the fittings are shown in Table 1. Inset: $T_{c}-T_{c}^{*}$ in log-log scale. Dashed lines are the corresponding power law fittings.
@@
We run simulations of the GH model defined on a Watts-Strogatz network of size $N$ with average degree $\langle k\rangle=12$ and rewiring probability $\pi=0.6$, from which we obtained both the average network activity $f_{a}$ and the susceptibility $\chi$ from Eq.(4) using both algorithms for $r_{1}=0$. The obtained susceptibility as a function of the threshold $T$ for different system sizes is shown in Fig.1 ($f_{a}$ as a function of $T$ will be analysed later, see Figs.6 and 7). We verified the presence of a size divergent peak of $\chi$. All the curves were fitted using appropriate fitting functions, from which the maxima of the curves were estimated as well as their location (pseudo critical points) $T_{c}^{*}(N)$. Using these values, we calculated the order parameter (activity) at the pseudo critical point $f_{a}(T=T_{c}^{*})$. The results for $T_{c}^{*}$ and $f_{a}(T=T_{c}^{*})$ as a function of $N$ are shown in Fig.2. We verified the expected finite size scaling behaviours in a critical phenomenon [15]: $\max\chi\sim N^{\gamma^{\prime}/\nu d}$, $T^{*}_{c}\sim T_{c}-N^{-1/\nu d}$ and $f_{a}(T=T^{*}_{c})\sim N^{-\beta/\nu d}$, where $\nu=\nu_{\perp}$ in conventional notation, and $d$ is the effective dimension. The estimated values of $T_{c}$ and the critical exponents are summarised in Table 1.
@@
Finally, the third method consists of running standard simulations for decreasing values of $r_{1}$, considering the $r_{1}\to 0$ limit. In Fig.3 we illustrate the typical behaviour of $\chi$ as a function of $T$ for different system sizes for three typical values of $r_{1}$. While for relatively large values of $r_{1}$ we recover the absence of scaling [4], as $r_{1}$ decreases, we see a clear emergence of a divergent peak with $N$. Following the same procedure as with the previous two methods, we estimated the critical exponents $\gamma^{\prime}/\nu d$, $\beta/\nu d$ and $1/\nu d$, for different values of $r_{1}$. The results are shown in Fig.4. We have a clear convergence when $r_{1}\to 0$ of all the exponents to values that are similar to the ones obtained by the two $r_{1}=0$ algorithms. The average values when $r_{1}\to 0$ are compared with those estimations in Table 1. We see that the three estimations are consistent, giving approximate values $\beta/\nu d\approx 0.3$ and $\gamma^{\prime}/\nu d\approx 0.25$ which, to the best of our knowledge, do not agree with any previously known universality class.
Critical phenomena on systems with an absorbing state exhibit the so-called generalized hyperscaling relation [11, 21]
$\gamma^{\prime}+\beta+\beta^{\prime}=\nu d$ (6)
where $\beta^{\prime}$ is the exponent related to the percolation probability [11]. In some systems (particularly those belonging to the Directed Percolation (DP) universality class) $\beta=\beta^{\prime}$ and Eq.(6) reduces to the usual hyperscaling relation $\gamma^{\prime}+2\beta=d\nu$ [11, 15]. This is due to the so-called rapidity reversal symmetry, which means that $f_{a}(t)\propto P_{sur}(t)$, where $P_{sur}(t)$ is the survival probability of finding at least one active site. In the Supplemental Material, we show evidence that, after a transient period, $f_{a}(t)\propto P_{sur}(t)$. In Table 1 we show the numerical value of $\gamma^{\prime}/(\nu d)+2\beta/(\nu d)$, obtained with the three different methods. These results suggest that $\gamma^{\prime}+2\beta<d\nu$, within the confidence intervals of our statistical error bars. However, systematic errors associated with the numerical methods, as well as possible difficulties to reach the large system size asymptotic regime, may introduce a bias, so the fulfilment of the hyperscaling relation $\gamma^{\prime}+2\beta=d\nu$ (as suggested by the supplemental material results) cannot be excluded.
Finally, we analysed the behaviour of the order parameter $f_{a}$ as a function of $r_{1}$ when $r_{1}\to 0$, for values of $T$ close to $T_{c}^{*}$. The numerical results are shown in the Supplemental Material. Our results are consistent with a power law behaviour $f_{a}\sim r_{1}^{1/\delta_{h}}$, not only at $T=T_{c}^{*}$, but also for a wide range of values of $T$ above $T_{c}^{*}$. The exponent $\delta_{h}\geq 1$ depends on $T$, converging to $\delta_{h}=1$ as $T$ increases to values far away from $T=T_{c}^{*}$. This behaviour suggests the presence of rare–regions effects [8, 20].
@@
Figure 3: Susceptibility as function of the threshold $T$ in Watts-Strogatz networks with $\langle k\rangle=12$ and $\pi=0.6$ for different network sizes NN and three typical values of the spontaneous activation probability: (a) $r_{1}=10^{-3}$, (b) $r_{1}=5\times 10^{-4}$ and (c) $r_{1}=10^{-5}$. The continuous lines are a guide to the eye.
@@ IV Activation mechanisms
Assuming that each neuron interacts with the average activity on the network, when $T$ is high enough, we can approximate $1-p_{n}(t)\simeq[1-a(t)P(W>T)]^{k}$, where 1a(t)P(W>T)1-a(t)P(W>T) is the probability of not being activated by a single active neighbour, and $k$ is the number of neighbour neurons (i.e., we do not consider here the possibility of a cooperative activation).
@@
Similarly, the equation accounts for random spontaneous activation and also for activation due to a single active presynaptic neuron (single activation, $a^{(s)}$). Cooperative activation due to the contribution of several presynaptic contributions $a^{(c)}$ is not considered there. In that sense, notice that the equation is quite similar to that derived in Ref.[7] for a different model which does not include cooperative activation at all. It is easy to see that Eq.(8) predicts a critical point at $r_{1}=0$ and $T_{c}=\ln(k)/\lambda$ where, close to the transition, we have $f_{a}(T,r_{1}=0)\sim(T-T_{c})^{\beta_{MF}}$, $\chi^{\prime}(T,r_{1}=0)={\partial f_{a}\over\partial r_{1}}|_{r_{1}=0}\sim(T-T_{c})^{-\gamma_{MF}}$ and $f_{a}(T=T_{c},r_{1})\sim r_{1}^{1/\delta_{h,MF}}$. The critical exponents are those of mean field directed percolation exponents: $\beta_{MF}=1$, $\gamma_{MF}=1$, and $\delta_{h,MF}=2$, in agreement with the results of Ref.[7]. We also see that, consistently with the results of section III, the spontaneous activation probability $r_{1}$ behaves as an external field conjugated to the order parameter $f_{a}$ of the continuous phase transition.
@@
Figure 7: Activities, susceptibility and $AC(1)$ as a function of $T$ for $r_{1}=10^{-5}$, $N=10000$, $\pi=0.6$ and different values of $\langle k\rangle$. Each column corresponds to one value of $\langle k\rangle$. From left to right $\langle k\rangle=2,8,12,40$. Upper row shows the activities as a function of $T$. Intermediate rows show the susceptibility $\chi$ and the first autocorrelation coefficient $AC(1)$. Blue dashed lines mark the pseudo critical thresholds $T_{c}^{*}$ obtained from $AC(1)$ peaks. The colour curves in the inset correspond to larger system sizes ($N=80000$ to 12800001280000) and show that the scaling peak of $\chi$ for $\langle k\rangle=8$ only emerges for large enough network sizes. The lower row shows the ratio among cooperative and single contributions to the activity, $f_{a}^{c}/f_{a}^{s}$ (spontaneous activation $f^{(sp)}=f_{a}-f_{a}^{(c)}-f_{a}^{(s)}$ is negligibly small and consequently not plotted). All simulations were averaged over at least 30 networks.
@@
In Fig.8 we compare the estimation of $T_{c}$ from the maximum of $AC(1)$ and the non- trivial (i.e., that scales with $N$) peak of the susceptibility as a function of $\langle k\rangle$ for $N=10000$ and $r_{1}=0$, using the reactivation algorithm (see section III). Checks were made for some values of $\langle k\rangle$ by extrapolating the $N\to\infty$ limit of $T_{c}^{*}$ in the continuous regime and also with standard simulations with $N=10000$ and $r_{1}=10^{-5}$. The differences were negligible in all the cases. We see that the estimations obtained from both quantities ($AC(1)$ and $\chi$) are indistinguishable for all values of $\langle k\rangle$. However the peak of $AC(1)$ is broad (i.e., it decays slowly as $|T-T_{c}|$ is increased) while the peak of $\chi$ may be narrow for some network topologies. Because of that, a reasonable estimate of $T_{c}$ can be found from $AC(1)$, for all studied values of $\langle k\rangle$, even for small values of $N$.